OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

Line perpendicular to the given line is X-Y+K=0 And put point to find k ie 2-(-2)+k=0 K=-4 so line is X-Y-4=0 Y=X-4

OpenStudy (anonymous):

life saver!!!!!

OpenStudy (anonymous):

ok